TIL UUIU WILTU U VUI LUIU 13 UISLILI UI LUMINOUS. TUIUU PUIG + 13. Classify probability. Rolling a fair 6-sided die. The prob please help solve #14.
Answer #1

Answer 14

The question can be modeled as a binomial distribution.

Let X be the number of correct answers. Then

X ~ Binomial(10,0.2)

P(X = 1) = (29) + 0.2+0,310-3

We need to calculate P(X>7) so that student passes

P(X > 7) = P(X = 7) + P(X = 8) + P(X = 9) + P(X = 10)

+ P(X > 7) = 0.0007864 + 0.0000737 +0.0000041 +0.0000001

+ P(X > 7) = 0.0008644

Calculations

X P(X=x)
0 0.1073742
1 0.2684355
2 0.3019899
3 0.2013266
4 0.0880804
5 0.0264241
6 0.0055050
7 0.0007864
8 0.0000737
9 0.0000041
10 0.0000001

Binomial Graph 0.3500000 0.3000000 0.2500000 0.2000000 Probability P(X=x) 0.1500000 0.1000000 0.0500000 0.0000000 9 10

Let me know in the comments if anything is not clear. I will reply ASAP! Please do upvote if satisfied!